OpenStudy (anonymous):

express f(x)= x^2-6x+14 in the form of f(x)= (x-h)^2=k

OpenStudy (anonymous):

express 14 as a sum of 9 and 5 and i guess it should be (x-h)^2-k

OpenStudy (anonymous):

I know what the actual answer is but I do not get how to get it. Is it like completing the square?

OpenStudy (anonymous):

you obtain an expressin of the form x^2-2xb+b^2 which can be written as (x-b)^2

OpenStudy (anonymous):

f(x) = x^2 - 6x + 14. You can find a perfect square in by taking (b/2)^2 = (-6/2)^2 = 9 f(x) = x^2 - 6x + (9 + 5) = (x-3)^2 + 5 (x-3)^2 = -5

OpenStudy (anonymous):

x^2-2*3*x+9+5=(x-3)^2+5

OpenStudy (anonymous):

Thank you!